OpenStudy (anonymous):

ax^2 + bx + c = 0, what happens when a = 0? What kind of graph will you get, and explain why?

OpenStudy (a_clan):

on putting a=0, we get bx+c=0 => x=-c/b which is the equation of a straight line. So, the graph is a vertical line intersecting the X-axis at x=-c/b

myininaya (myininaya):

y=ax^2+bx+c of a=0 then y=bx+c which is a line